a block of mass m =2 kg is pulled by a force f=40 N upward through a height h =2 m . find the work done on the block by the applied force F and its weight mg

Grade:9

1 Answers

Arun
6 years ago
Mass of block = 2Kg 

F = 40 N 

h = 2m 

Work done by the force F = F*h = 40*2= 80 Joule and, 

Work done by weight mg = -mgh = -2*10*2 = -40 J